OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {14, 18, 21, 15, 17} Data Set 2: {15, 17, 22, 20, 16}

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@robtobey @texaschic101 @dan815

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@Jaynator495 @JoannaBlackwelder

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

I could've swore we did this one today, we have to find the mean values of both data sets, first.. the first one is 83. the second one is 76......now, we have to compute the MAD of each data set. THE MAD would equal: MAD=26/5=5.2, However then the ratio mean-to.-MAD, for first data set, is: 83/5.2=...? (15.96153846) 15.96. Comparing wise: 75−76=1 74−76=2 68−76=8 83−76=7 80−76=4. ============ we have: -1, -2, -8, 7, 4 ------- So in conclusion our first line would be: 15.96 Then our second line would be: 16.0 Hope this helps, have a nice day. Tag me on anymore you need help with (:

OpenStudy (anonymous):

@Hotchellerae21